OpenStudy (anonymous):

The perimeter of a rectangular store receipt is 10 5/6 inches. If three of the sides total 8 1/2 inches, what is the length of the fourth side?

OpenStudy (anonymous):

@omnomnom

OpenStudy (anonymous):

OMNOMNOMNOMNOMNOMNOMNOMNOMNOMNOMNOMNOMNOMN. That was some tasty burrito

OpenStudy (omnomnom):

xD lol what burrito?

OpenStudy (omnomnom):

\[10 \frac{ 5 }{ 6 }\] \[8 \frac{ 1 }{ 2 }\] These are the numbers right?

OpenStudy (anonymous):

no

OpenStudy (omnomnom):

then what are they? -_-

OpenStudy (anonymous):

they are numbers.

OpenStudy (omnomnom):

i know that but.. omg how do i explain..... are they fractions?

OpenStudy (anonymous):

no they are decimals.

OpenStudy (omnomnom):

then write them like decimals man -_-

OpenStudy (anonymous):

no they are people treat them with respect.

OpenStudy (omnomnom):

O_O Lol xD

OpenStudy (anonymous):

So do you know the answer wanna help me lol.

OpenStudy (omnomnom):

well if you write the numbers i will cuz i cant understand the numbers.. -_- they look like fractions to me

OpenStudy (anonymous):

They are

OpenStudy (omnomnom):

then why did you say decimals? T_T

OpenStudy (omnomnom):

first of all what grade are you in? -_- this is a easy question.. or at least it looks like one

OpenStudy (omnomnom):

subtract the fractions after finding the same denominator

OpenStudy (anonymous):

is it 3

OpenStudy (omnomnom):

hmm..... hold on just a sec let moi see :P

OpenStudy (omnomnom):

No thats not the answer the answer is \[2 \frac{ 1 }{ 3 }\]

OpenStudy (omnomnom):

Ello ? :O

OpenStudy (omnomnom):

I MIGHT BE WRONG SO DONT USE MY ANSWER